# Functional Brain-to-Brain Transformation with No Shared Data

### v1

@@ 3.2 Improving Image-to-fMRI Encoding of one Subject using another Higher-Quality Subject
Figure 6: Improving encoding of a subject from Vim1 3-Tesla dataset (“student”) using another high-quality subject (“teacher”): This figure compares the quality of the “student” baseline encoder (which is trained solely on the student’s own fMRI data), to the student encoder obtained when incorporating also data from the “teacher” during training. In each plot, the grey line represents the student baseline encoder model, the purple line symbolizes the encoder improvement derived using as a “teacher” from the GOD dataset (e.g., Subject4 is the highest-quality subject in the GOD dataset). The orange line corresponds to the encoder improvement achieved by utilizing a superior subject from another dataset (NSD), which has more scanned examples and higher fMRI resolution (7-Tesla machine). The quality of the resulting encoders is assessed through the mean Pearson correlation of all predicted fMRI voxels. The “number of training examples” refers to the number of the student’s own image-fMRI pairs of examples used for training the “student” encoder, whereas the “teacher” subject (drawn from either GOD or NSD dataset) has been trained using all its own available examples and remains unchanged during the training of the student encoder. (a) Results for an individual student subject (Subject1). (b) Results for an individual student subject (Subject2). (c) The distribution of Pearson correlations across the fMRI voxels, comparing the baseline student encoder against the improved student encoder when trained with a teacher from the NSD dataset. This is evaluated using varying numbers of training examples (300, 900, and 1750) from the student’s own fMRI data. The x-axis signifies the Pearson correlation values, while the y-axis represents the density of fMRI voxels corresponding to each correlation value.
Figure 5 compares the quality of the “student” baseline encoder (which is trained solely on the student’s own fMRI data), to the student encoder obtained when incorporating also data from the “teacher” during training. The student encoder benefits from the unique fMRI data of the teacher (the subject with higher-quality data, e.g., from NSD 7 Tesla dataset), as well as from images that neither subjects have encountered (via the “teacher” higher quality encoder). This data is used in addition to the student’s own fMRI data. Improvements of subjects’ encoders from the GOD 3-Tesla dataset are depicted in Figure 5. In each plot, the grey line represents the student baseline encoder model, the purple line symbolizes the encoder improvement obtained when using a “teacher” from the same dataset (e.g., Subject4 is the highest-quality subject in the GOD dataset). The orange line corresponds to the encoder improvement achieved by utilizing a superior subject from another dataset (NSD), which has more scanned examples and higher fMRI resolution (7-Tesla machine).
The quality of the resulting encoders is assessed through the mean Pearson correlation of all predicted fMRI voxels. The “number of training examples” refers to the number of the student’s own image-fMRI pairs of examples used for training the “student” encoder (Subject1 from GOD dataset), whereas the “teacher” encoder (of a different subject from either GOD or NSD dataset) has been trained using all its own available examples and remains unchanged during the training of the student encoder.
@@
Figure 5.a shows the results for an individual subject (Subject1), and Figure 5.b aggregates the results of all five subjects in GOD dataset (showing mean values and SEM). Both graphs show that encoders trained with a “teacher” subject exhibit significant improvement over their baseline encoder. Notably, employing Subject4 from the same dataset as a “teacher” results in a substantial improvement of approximately 50% ($df=20$, $t=8.15$, $p<5e-7$, $d=1.82$). Remarkably, utilizing a subject from the higher-resolution NSD dataset as a “teacher” leads to even more pronounced performance gains. Specifically, for a small number of examples, it demonstrates a nearly 100% relative improvement over the baseline encoder, and 50% over the other enhanced encoder model ($df=20$, $t=5.85$, $p<5e-5$, $d=1.31$). These observations are remarkable given that the NSD dataset does not share images with the GOD dataset, whereas a “teacher” subject from the same dataset shares numerous images with the “student” subject. Supplementary examples showcasing the improvement of specific subjects’ encoders are provided in Figure LABEL:S6 in the Supplementary Material.
Lastly, Figure 5.c presents the distribution of Pearson correlations across the fMRI voxels, comparing the baseline student encoder against the improved student encoder when trained with a teacher from the NSD dataset. This is evaluated using varying numbers of training examples (300, 600, and 1200) from the student’s own fMRI data. As the figure indicates, our enhanced encoder (depicted by the orange histogram) shifts the distribution to the right, signifying the improved correlation in a substantial portion of fMRI voxels.
@@
Next, we present our results on the old Vim-1 fMRI dataset, which contains only two subjects who observed grey-scale images. Given the limited subject number, we showcase the method using a “teacher” subject drawn from other datasets, namely GOD and NSD. Figure 6.a and 6.b shows the results for each individual subject. As in Figure 5, the grey line represents the student baseline encoder, the purple line symbolizes the encoder improvement obtained using as a “teacher” from the GOD dataset. The orange line corresponds to the encoder improvement achieved by utilizing a superior subject from the higher fMRI resolution dataset (NSD). Both plots show that encoders trained with a “teacher” subject exhibit an improvement over their baseline encoder model. The results illustrate that, as expected, a subject from the NSD 7-Tesla dataset, when used as a “teacher”, outperforms using a “teacher” subject from the GOD 3-Tesla dataset. Notably, despite the considerable differences between the “teacher” and “student” subjects in terms of data characteristics (color vs grey-scale images, different image datasets), the method remains effective in leveraging one subject’s higher-quality data to improve another.
Lastly, Figure 6.c presents the distribution of Pearson correlations across the fMRI voxels, comparing the baseline student encoder against the improved student encoder when trained with a teacher from the NSD dataset. This is evaluated using varying numbers of training examples (300, 900, and 1750) from the student’s own fMRI data. As the figure indicates, our enhanced encoder (depicted by the orange histogram) shifts the distribution to the right, signifying the improved correlation in a substantial portion of fMRI voxels.
We further examined the usefulness of our improved encoders (compared to baseline encoders) for “Zero-Shot” classification (i.e., classification of fMRI data to novel image categories, which were never-before-seen during training). The classification method we used is similar to that of [30] (with some small modifications), and is described in the supplementary text and in Figure LABEL:S7. Figure 7 compare the classification performance using different encoders. The results indicate that, irrespective of the number of training examples, the enhanced encoders (represented by the purple column) achieve significantly higher classification accuracy than the baseline encoders (grey column), as confirmed by statistical analysis (paired samples student’s t-test, $df=20$, $t=4.69$, $p<5e-5$, $d=1.05$). Furthermore, the application of an enhanced encoder trained with the 7-Tesla dataset (orange column) demonstrates even greater classification accuracy, supported by a t-test comparison with the GOD 3-Tesla “teacher” ($df=20$, $t=3.76$, $p<5e-3$, $d=0.84$). Additionally, Figure LABEL:S8 includes results showing the improved brain-to-brain transformations within the GOD dataset capability of the improved encoder to incorporate “external” images with shared data, further underscoring their effectiveness in enhancing various tasks in lower-resolution datasets.
Figure 7: Improving classification accuracy (of never-before-seen image classes) of a 3-T GOD subject, using its enhanced encoder via another subject: This plot presents the classification results employing this approach, considering different modes of encoder training. We compare the baseline encoder (trained only on its subject-specific fMRI data) with encoders improved by another subject (“teacher”), either from the GOD dataset or using the NSD 7-Tesla dataset. The results presented are the mean accuracy of all subjects with the SEM error.
@@ 4 Discussion
This paper introduces, for the first time, an approach to learn functional brain-to-brain transformations between different subjects without requiring any shared data. Our methods hold the potential to significantly enhance research in brain functionality, expanding both the quantity and quality of data available for analysis. This breakthrough offers exciting research opportunities, enabling the merging and enrichment of diverse datasets, even when they involve different stimuli or were collected using fMRI machines of varying resolutions.
Beyond the novel method for computing brain transformations without shared data, our innovation also lies in the integration of visual encoders with these transformations. We can now utilize both “non-shared” and “external” images for training encoders and transformations. This ability to incorporate images unseen by any of the subjects marks a significant advancement, greatly expanding the capacity of training data for various different fMRI related tasks.
Our method is particularly effective in utilizing new high-quality fMRI datasets, such as the NSD 7-Tesla dataset, to enhance older, lower-quality datasets like the GOD and Vim1 3-Tesla datasets. We have demonstrated substantial improvements in image-to-fMRI encoding with low-resolution 3T datasets. These enhanced encoders further enable improvements in within-dataset brain-to-brain transformations and boost the accuracy of semantic classification for novel, never-before-seen image categories. These enhancements are particularly exciting, offering the potential to significantly advance our understanding of how the brain perceives and processes visual information. Altogether, this underscores the importance of our method in improving older datasets by leveraging newer, high-quality datasets.
Importantly, our approach is general, hence can be extended to other modalities (e.g., MEG, EEG, ECoG), as well as to other stimulus types (e.g., audio, video, text). We further believe that our proposed approach has numerous potential applications, which may advance our understanding of the brain and its functions.
@@ 5.1 Datasets
We tested our approach on three distinct publicly available fMRI datasets: “Natural Scenes Data set“ (NSD) [24], Generic Object Decoding (GOD) [23], and “Visual Imaging 1” (Vim1) [31, 33]. The datasets provide fMRI responses of human subjects to a variety of natural images. In all the experiments, the participants were instructed to fixate their eyes on the center of the presented images. We summarize differences between the datasets in Table 1.
@@
Natural Scenes Dataset (NSD) is a new dataset recorded using 7-Tesla fMRI machine, resulting in higher voxel resolution. All 8 subjects in this dataset were presented with 1,000 shared images and 9,000 unique images per subject. This is contrary to other datasets where all participants were presented the same stimuli. Vim1 dataset has grey scale stimuli images with a circular mask (see Figure 3), whereas the other datasets that have RGB images. It should be noted that each dataset was scanned with a different machine as well as different recording scheme and processing protocols. We use the same approach and architecture to produce results for all the datasets. In addition to the fMRI datasets, we also used 50,000 natural images as “External Images” (images without any fMRI). Those images were taken from 1000 classes of ImageNet validation data (“ILSVRC” [38]). Those images are the “external” data reported in our experiments.
@@
Dataset Number of distinct stimuli Number of subjects Machine resolution Image origin
NSD [24] 10000 per subject (7̃3000 total) 8 7T COCO  [39]
GOD [23] 1250 per subject (1̃250 total) 5 3T ImageNet  [38]
Vim1 [31, 33] 2000 per subject (2̃000 total) 2 3T Multiple sources
Table 1: Summary of fMRI datasets used in the paper.
5.2 Technical Details of the Method
5.2.1 Training brain-to-brain transformations with no shared data

### v2

@@ 3 Results
Figure 6: Improving Image-to-fMRI Encoding via B2B Transformation: This figure compares the quality of the “student” baseline encoder (which is trained solely on the student’s own fMRI data) to the student encoder obtained when incorporating also data from the “teacher” via our B2B transformation). In each plot, the grey line represents the student baseline encoder model, the purple line represents the encoder improvement obtained when using a “teacher” from the GOD dataset (e.g., Subject4 is the highest-quality subject in the GOD dataset). The orange line corresponds to the encoder improvement achieved by utilizing a superior subject from another dataset (NSD), which has more scanned examples and higher fMRI resolution (7-Tesla machine). The quality of the resulting encoders is assessed through the mean Pearson correlation of all predicted fMRI voxels. The “number of training examples” refers to the number of the student’s own image-fMRI pairs of examples used for training the “student” encoder, whereas the “teacher” subject (drawn from either GOD or NSD dataset) has been trained using all its own available examples and remains unchanged during the training of the student encoder. The dotted lines mark which two types of models were statistically compared, and the number of asterisks (*) mark the significance (see Methods). (a) Mean Pearson correlation results averaged over all five subjects in the GOD dataset, as well as individual subject results. (b) Mean Pearson correlation results averaged across the two subjects in the Vim1 dataset, as well as individual subject results.
Figure 7: Image Retrieval via encoded fMRI: (a) Image retrieval accuracy based on encoded fMRI, showing top-1 accuracy by comparing the test fMRI to 1000 encoded fMRIs, averaged across all subjects and for individual subjects. The image retrieval via fMRI predicted by a B2B-enhanced encoder (Orange line – with a strong NSD teacher), is typically 200%-500% more accurate than retrieval from an fMRI predicted by the Baseline encoder (Grey line – same encoder, but without a teacher). (b) The figure shows the top 3 retrieved images for each test fMRI, comparing the baseline encoder with the enhanced encoder trained using an NSD teacher. For each test case, 1000 candidate images were used, and the top 3 matches are presented.
In all the results presented next, grey color represents the student baseline encoder (trained solely on the student’s fMRI data), purple color illustrates the improvement using a teacher from the same dataset (GOD), and orange color demonstrates the greater improvement achieved with a teacher from the 7-Tesla dataset (NSD). The “number of training examples” refers to the student’s own image-fMRI pairs, while the teacher encoder (trained on all available examples) remains unchanged during the student’s training process.
@@
(i) Pearson Correlation Evaluation: Figure 6.a shows mean Pearson correlation results for individual subjects (Subject2 and Subject3), and averaged over all five subjects in GOD dataset (showing mean values and SEM on GOD test set). Those graphs show that Image-to-fMRI encoders trained with a “teacher” subject (via our B2B transformation), exhibit a significant improvement over their baseline encoder (trained only on the “student”’s data). Notably, employing Subject4 from the same dataset as a “teacher” results in a substantial improvement of approximately 50% over the baseline encoder ($N=48$, $p<0.00001$, $d=4.703$). Remarkably, utilizing a subject from the higher-resolution NSD dataset as a “teacher” leads to even more pronounced performance gains. Specifically, for a small number of examples, it demonstrates a nearly 100% relative improvement over the baseline encoder, and 10%-30% over the other enhanced encoder model ($N=48$, $p<0.0001$, $d=3.143$). These observations are remarkable given that the NSD dataset does not share any images with the GOD dataset and was scanned on a different fMRI machine, whereas a “teacher” subject from the same dataset shares numerous images with the “student” subject (and was scanned on the same fMRI machine). Supplementary examples showcasing the improvement of all specific subjects’ encoders are provided in Figure S9 in the Supplementary Material.
@@
Figure 6.b further presents results of enhancing Image-to-fMRI encoding on the old Vim-1 fMRI dataset. This dataset contains only two subjects who viewed grey-scale images. Given the small number of subjects, we showcase the method using a “teacher” subject from other datasets, namely GOD or NSD. The plots show that encoders trained with a “teacher” subject (whether from GOD or NSD) exhibit an improvement over their baseline Vim-1 encoder model. Moreover, as expected, a “teacher” from the NSD 7-Tesla dataset outperforms a “teacher” from the GOD 3-Tesla dataset. Notably, despite the considerable differences between the “teacher” and “student” subjects in terms of data characteristics (Color vs Grey-scale images, and different fMRI machines), the method remains effective in leveraging one subject’s higher-quality data to improve another.
(ii) Image Retrieval via Encoded fMRI: We further assess the impact of our B2B-enhanced encoding through an fMRI retrieval test. For each fMRI scan in the test set (denoted as “Query”), we aim to retrieve (detect) its corresponding Test-image which produced it out of a set of N images (the Test-image and N-1 random distractors). To do so, we first predict the fMRIs of all N images in the set (using either the Baseline encoder or the B2B-enhanced encoder). We then search for the Nearest-Neighbor (NN) of the real test-fMRI (Query) among the set of N predicted fMRIs (using cosine similarity). If the fMRI predicted from the Test-image is retrieved as the $1^{st}$ NN, it obtains a “Rank-1” score. If it is retrieved as the $k^{th}$ NN, it obtains a “Rank-k” score. The reported Top-1 accuracy is the percent of test-fMRIs which obtained a Rank-1 score.
Figure 7.a shows significant improvement in retrieval accuracy of GOD “student” encoders, when enhanced with a “teacher” either from the GOD dataset or from the NSD dataset. Each Test-fMRI is compared against 1000 encoded fMRIs – from the real Test-image, and from 999 other ‘distractor’ images from ImageNet. The B2B-enhanced encoder, especially when using an NSD teacher, outperforms the Baseline encoder. Please note that image retrieval via fMRI predicted by a B2B-enhanced encoder (Orange line – with a strong NSD teacher), is typically 200%-500% more accurate than retrieval from an fMRI predicted by the Baseline encoder (Grey line – same encoder, but without a teacher). This shows the significant power of our B2B transformations across brains and datasets, despite having no shared data.
Figure 7.b visually presents qualitative retrieval results comparing the Baseline Encoder to the B2B-Enhanced Encoder (trained with the NSD “teacher”). For each test fMRI, we show the top 3 retrieved images (rank 1, 2, and 3) out of 1000. As seen, the B2B-Enhanced encoder not only retrieves the correct image more accurately (lower rank), but also the $2^{nd}$ and $3^{rd}$ ranked images are often from the same semantic category or visually similar to the correct image. This suggests that the B2B-Enhanced encoder generates fMRIs which capture more nuanced similarities, keeping related images closer in fMRI space. This further demonstrates its effectiveness in better representing brain responses.
(iii) Image Classification from fMRI: We further evaluated the significance of our enhanced fMRI encoding for “Zero-Shot” classification – namely, classification of fMRI data into new image categories that were never seen during training. We devised a simple fMRI-classification algorithm, which uses an Image-to-fMRI encoder to predict fMRI patterns for 100 arbitrary images per class (since these are never-before seen image categories for which there are no real fMRI training examples). These 100 predicted fMRIs are then averaged to form an “fMRI Representative” per category. In the classification stage, each test-fMRI (a real fMRI scan) is compared against these generated fMRI class representatives using cosine similarity. The most similar fMRI-representative determines the category classification. In an $n$-way classification task, the test-fMRI is compared against $n$ fMRI class-representatives: the generated fMRI-representative of the correct image class, along with $n$-1 generated fMRI-representatives of other $n$-1 randomly selected ImageNet classes.
Figure 8 compares the $100$-way classification accuracy (out of $n$=100 classes) using Baseline vs. B2B-enhanced encoders. The results indicate that, irrespective of the number of training examples, the B2B-enhanced encoders (represented by the purple column) achieve significantly higher classification accuracy (paired samples permutation test, $N=20$, $p<0.00001$, $d=1.01$) than the baseline encoders (grey column). Furthermore, the B2B-enhanced encoder trained with the 7-Tesla “teacher” (orange column) demonstrates even greater classification accuracy than the GOD 3-Tesla “teacher” ($N=20$, $p<0.00001$, $d=0.98$). In particular, note that image classification from fMRI predicted by a B2B-enhanced encoder (Orange bar – with a strong NSD teacher), is typically $\sim$20%-50more accurate than image classification from an fMRI predicted by the Baseline encoder (Grey bar – same encoder, but without a teacher), and 1500%-2000% more accurate than chance level (Grey dashed-line). Recall that all $n$=100 image classes/categories were never seen during training time, neither by the teacher, nor by the baseline encoder.
Figure 8: Image Classification from fMRI of 3-T GOD subjects: This plot presents the classification accuracy (out of 100 classes) using Baseline vs. B2B-enhanced encoder. The results presented are the mean accuracy of all subjects with the SEM error and individual subject result with mean and SEM over 3 training repetitions. The image classification from fMRI predicted by a B2B-enhanced encoder (Orange bar – with a strong NSD teacher), is typically $\sim$20%-50more accurate than image classification from an fMRI predicted by the Baseline encoder (Grey bar – same encoder, but without a teacher), and 1500%-2000% more accurate than chance level (Grey dashed-line). The $n$=100 image classes/categories were never seen during training time, neither by the teacher, nor by the baseline encoder..
Analysis and ablations of the various factors which contribute to the improved encoder performance are provided in the Supplementary Section S1. In particular, it analyzes the influence of data quality and quantity, showing that both factors contribute to NSD’s superiority as a dataset for enhancing other subjects’ encoders.
@@ 4 Discussion
This paper introduces, for the first time, an approach to learn functional brain-to-brain (B2B) transformations between different subjects without requiring any shared data. Our approach enables the merging and enrichment of diverse fMRI datasets collected around the globe, even when they were collected on different image stimuli and on different fMRI machines of varying resolutions. This may provide a significant boost in brain research, by increasing both the quantity and quality of fMRI data available for analysis. We show that it is possible to utilize both “non-shared” and “external” images to train functional B2B transformations, and to significantly improve Image-to-fMRI encoders. This ability to incorporate images unseen by any of the subjects marks a significant advancement, greatly expanding the number of available training data (by several orders of magnitude). Our method is particularly effective in utilizing new high-quality fMRI datasets, such as the NSD 7-Tesla dataset, to enhance Image-to-fMRI encoders of subjects in older, lower-quality datasets like the GOD (3-Tesla) and Vim1 (4-Tesla) datasets. We have demonstrated that these B2B-Enhanced Encoders further improve the accuracy of “zero-shot” semantic image classification directly from fMRI, as well as provide improved fMRI retrieval accuracy.
Our method was so far demonstrated effectively on fMRI data in response to visual stimuli. Extending our approach to to other sensing modalities (e.g., MEG, EEG, ECoG), or to other types of data stimuli (e.g., audio, video, text) has great potential. As with any Deep-Learning based method, the performance of our approach may deteriorate when applied to out-of-distribution images that differ significantly from the training set. Therefore, both our B2B transformation and enhanced encoder may have limited applicability when datasets differ significantly in image characteristics. Despite this, there are likely underlying similarities in image features and corresponding brain activations that our method can leverage, although with potentially reduced effectiveness. Finally, the quality of the encoder is a key factor in our method’s success. While small datasets with limited number of samples per subject can limit the overall quality of the results, our findings indicate that our method can still have good prediction correlation with as few as 300 image-fMRI example pairs.
Our capability to aggregate data across diverse datasets and conditions without requiring shared data, alongside improvements in encoder performance, provides unique opportunities for advancing cognitive neuroscience and computational neuroimaging. By facilitating direct comparisons between subjects who viewed different image sets, recorded on various machines in labs worldwide, our B2B-transformation approach opens the door to broader, cross-dataset analyses which were previously unattainable, with many new potential applications. For example, our B2B-transformations may facilitate the integration of fMRI datasets of different behavioral tasks across different fMRI datasets. It may create new opportunities to study a range of visual and cognitive tasks collectively — a level of analysis challenging, if not impossible, without a mechanism to bridge data collected under varied conditions. Our approach may also support future studies of brain functionality across individuals with diverse backgrounds or health conditions, whether they are viewing images or, potentially, engaging with other stimuli types like video and audio.
Our B2B-enhanced Image-to-fMRI encoders also open the door to many new potential opportunities. Their effective performance with minimal data may be particularly valuable for labs that cannot collect extensive datasets but aim to investigate specific scenarios. Due to the encoder’s ability to predict fMRI activations on a vast number of images, researchers may explore and compare fMRI responses to a wider variety of stimuli. For example, compare fMRI responses to images of indoor versus outdoor scenes, or even contrast fMRI responses to emotional content (e.g., happy versus sad images). All in all, our methods for training B2B transformations without shared data and using those to enhance Image-to-fMRI encoding could serve as valuable tools for future research.
@@ 5.1 Datasets
We tested our approach on three distinct publicly available fMRI datasets: “Natural Scenes Data set“ (NSD) [26], Generic Object Decoding (GOD) [25], and “Visual Imaging 1” (Vim1) [33, 35]. The datasets provide fMRI responses of human subjects to a variety of natural images. We summarize differences between the datasets in the table below.
@@
Natural Scenes Dataset (NSD) is a new dataset recorded using 7-Tesla fMRI machine, resulting in higher voxel resolution. All 8 subjects in this dataset were presented with 1,000 shared images and 9,000 unique images per subject. This is contrary to other datasets where all participants were presented the same stimuli. Vim1 dataset has grey scale stimuli images with a circular mask (see Figure 3), whereas the other datasets that have RGB images. It should be noted that each dataset was scanned with a different machine as well as different recording scheme and processing protocols. We use the same approach and architecture to produce results for all the datasets. In addition to the fMRI datasets, we also used 50,000 natural images as “External data” (images without any fMRI). Those images were taken from 1000 classes of ImageNet validation data (“ILSVRC” [40]). Those images are the “external” data reported in our experiments. For both the VIM-1 and GOD datasets, we utilized the predefined test sets, which consist of 120 images for VIM-1 and 50 images for GOD. We reserved 10% of the training set to serve as a validation set. In the case of NSD, since there was no dedicated test set, we allocated 10% of the data for validation and another 10% for testing. Both validation and test sets were proportionally sampled from the shared data across all subjects and the non-shared data. This approach was crucial to ensure that the test set included shared data, enabling the evaluation of transformations that can only be assessed using shared data across subjects.
@@
Dataset Number of distinct stimuli Subjects Voxels Resolution Image origin
NSD [26] 10000 per subject ($\sim{73000}$ total) 8 10000 7T COCO  [41]
GOD [25] 1250 per subject (1250 total) 5 5000 3T ImageNet  [40]
Vim1 [33, 35] 2000 per subject (2000 total) 2 5000 4T Multiple sources
5.2 Data Acquisition and Processing
The datasets utilized in our study include BOLD fMRI responses to various natural images, recorded over multiple scanning sessions. In all experiments, participants were instructed to fixate their eyes on the center of the presented images. Each dataset underwent specific pre-processing procedures as detailed in their respective publications [33, 35, 25, 26]. Building on these procedures, we implemented additional processing steps:
Voxel Selection.
Across all datasets, a subset of brain voxels relevant to visual stimuli was selected. For the GOD dataset, we utilized a predefined selection of 5,000 voxels from the visual areas. In the NSD and VIM-1 datasets, we applied a data-driven approach to voxel selection based on the signal-to-noise ratio (SNR). SNR was computed as the ratio of the variance in voxel responses to different stimuli (reflecting true neural activity) to the variance in responses to the same stimulus across multiple repetitions (reflecting noise). In NSD, each image had between one and three repetitions, and in VIM-1, each image had two repetitions in the training set. By selecting the top voxels with the highest SNR, we ensured that the chosen voxels were those that most reliably reflected neural responses to visual stimuli. Specifically, we selected the top 5,000 voxels for VIM-1 to align with the GOD dataset, and the top 10,000 voxels for NSD , reflecting its higher resolution and the greater number of scanned images. In both dataset voxels were chosen from the entire brain.
Voxel Z-scoring.
Each voxel underwent Z-scoring normalization within each run, defined as a continuous fMRI scanning period. This normalization standardized voxel responses across runs, improving data comparability and consistency.
Averaging across repetitions.
To enhance fMRI data quality, images were presented multiple times to each subject, allowing for higher-quality measurements. In VIM-1, each training image was repeated twice, and each test image 13 times. NSD images had between one and three repetitions. We averaged the fMRI responses across repetitions to improve signal quality. For the GOD dataset, we used a version with five repetitions per image in the training set and 24 in the test set. To illustrate the impact of repetition, we present in the supplementary results using both single-repetition fMRI and averaged fMRI with varying repetition counts.
Group space fMRI - fsaverage.
For comparative analysis between anatomical and functional alignments, we used the ’fsaverage’ shared space provided by FreeSurfer [12, 13] only for the transformation evaluation in the NSD dataset. This approach enabled a fair and direct comparison by performing both alignments in a common reference space. Specifically, we selected a consistent set of 10,000 voxels across all subjects, based on the SNR of Subject 1. The fsaverage space was derived using pre-calculated betas from the NSD This procedure involved resampling subject-native cortical surfaces at three different depths using cubic interpolation. Subsequently, the averaged betas were produced by averaging these resampled values. Finally, these averaged betas were mapped onto the fsaverage surface through nearest-neighbor interpolation. This methodical approach allows us to compare functional alignment results directly with those obtained from anatomical alignment, each within the same standardized brain space.
5.3 Visual Encoder
The image-to-fMRI encoder employed in our work is based on the framework outlined in [32]. It encompasses a pre-trained VGG network (trained on ImageNet classification task), augmented with additional convolution layers and a non-linear layer to project data into the fMRI space. First, the input image is passed through VGG19 network [42], intermediate embeddings of the network blocks 1-4 are extracted, each compromising a parallel branch of feature representation. Each of the branches feature map is reduced spatially and the channels’ dimension as well to 28$\times$28$\times$32 (Height×$\times$×Width×$\times$×ConvolutionChannels). This is done by a series of modules consisting of 3$\times$3 convolution with 32 channels, ReLU, $\times$2 sub-sampling, and batch normalization. On the reduced feature maps, we train a spatial layer that transfers the spatial position of the feature map to voxels. For each voxel, we learn a linear combination of the spatial positions (where channels are not affected). This is done for each one of the branches and the layer weights are shared between the branches. After this operation we have a representation with dimensions Voxels×$\times$×ConvolutionChannels. This is followed by a locally connected layer with output size of 1, this layer learns how to combine the channels for each one of the voxels, representation with dimensions Voxels×$\times$1, for each one of the branches. The final layer combines the results of the 4 branches, this is done by concatenating the 4 branches and applying a locally connected layer with positive weights (averaging).
5.4 Training brain-to-brain transformations with no shared data